I don't understand solving equations. The problem is:
3x+2y= -9
-10x+5y= -5

Answer by mananth(16946)

3x+2y=-9 .............1
-10x+5y =-5 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 2
-15x-10y=45
-20x+10 y=-10
Add the two equations
-35x=35
/ -35
x = -1
plug value of x in (1)
3 x + 2 y = -9
-3 + 2 y = -9
2 y = -9 + 3
2 y = -6
y = -3